\begin{thebibliography}{99}

\bibitem{Hua-Qi-toader312}
H. Alzer and S.-L. Qiu, \emph{Monotonicity theorems and inequalities for the complete elliptic integrals}, J. Comput. Appl. Math. \textbf{172} (2004), no.~2, 289\nobreakdash--312; Available online at \url{http://dx.doi.org/10.1016/j.cam.2004.02.009}.

\bibitem{Hua-Qi-toader38}
G. D. Anderson, M. K. Vamanamurthy, and M. Vuorinen, \emph{Conformal Invariants, Inequalities, and Quasiconformal Maps}, John Wiley \& Sons, New York, 1997.

\bibitem{Hua-Qi-toader311}
R. W. Barnard, K. Pearce, and K. C. Richards, \emph{An inequality involving the generalized hypergeometric function and the arc length of an ellipse}, SIAM J. Math. Anal. \textbf{31} (2000), no.~3, 693\nobreakdash--699; Available online at \url{http://dx.doi.org/10.1137/S0036141098341575}.

\bibitem{Hua-Qi-toader36}
F. Bowman, \emph{Introduction to Elliptic Functions with Applications}, Dover Publications, New York, 1961.

\bibitem{Hua-Qi-toader37}
P. F. Byrd and M. D. Friedman, \emph{Handbook of Elliptic Integrals for Engineers and Scientists}, Springer-Verlag, New York, 1971.

\bibitem{Hua-Qi-toader35}
Y.-M. Chu and M.-K. Wang, \emph{Inequalities between arithmetic-geometric, Gini, and Toader means}, Abstr. Appl. Anal. \textbf{2012}, Article ID 830585, 11 pages; Avaiable online at \url{http://dx.doi.org/10.1155/2012/830585}.

\bibitem{Hua-Qi-toader33}
Y.-M. Chu and M.-K. Wang, \emph{Optimal Lehmer mean bounds for the Toader mean}, Results Math. \textbf{61} (2012), no.~3-4, 223\nobreakdash--229; Available online at \url{http://dx.doi.org/10.1007/s00025-010-0090-9}.

\bibitem{Hua-Qi-toader313}
Y.-M. Chu, M.-K. Wang, and X.-Y. Ma, \emph{Sharp bounds for Toader mean in terms of contraharmonic mean with applications}, J. Math. Inequal. \textbf{7} (2013), no.~2, 161\nobreakdash--166; Available online at \url{http://dx.doi.org/10.7153/jmi-07-15}.

\bibitem{Hua-Qi-toader34}
Y.-M. Chu, M.-K. Wang, and S.-L. Qiu, \emph{Optimal combinations bounds of root-square and arithmetic means for Toader mean}, Proc. Indian Acad. Sci. Math. Sci. \textbf{122} (2012), no.~1, 41\nobreakdash--51; Available online at \url{http://dx.doi.org/10.1007/s12044-012-0062-y}.

\bibitem{Jiang-toader2.tex}
Y. Hua and F. Qi, \textit{The best bounds for Toader mean in terms of the centroidal and arithmetic means}, available online at \url{http://arxiv.org/abs/1303.2451}.

\bibitem{jiang-qi-toader-bound.tex}
W.-D. Jiang and F. Qi, \textit{Bounds for the combination of Toader mean and the arithmetic mean in terms of the contraharmonic mean}, available online at \url{http://arxiv.org/abs/1402.4561}.

\bibitem{Jiang-Neuman.tex}
W.-D. Jiang and F. Qi, \textit{Sharp bounds for Neuman-S\'andor's mean in terms of the root-mean-square}, available online at \url{http://arxiv.org/abs/1301.3267}.

\bibitem{jiang-qi-October.tex}
W.-D. Jiang and F. Qi, \textit{Sharp bounds in terms of the power of the contra-harmonic mean for Neuman-S\'andor mean}, available online at \url{http://arxiv.org/abs/1301.3554}.

\bibitem{background-Jiang-Qi.tex}
W.-D. Jiang and F. Qi, \textit{Some sharp inequalities involving Seiffert and other means and their concise proofs}, available online at \url{http://arxiv.org/abs/1201.6480}.

\bibitem{background-Jiang-Qi-revised.tex}
W.-D. Jiang and F. Qi, \emph{Some sharp inequalities involving Seiffert and other means and their concise proofs}, Math. Inequal. Appl. \textbf{15} (2012), no.~4, 1007\nobreakdash--1017; Available online at \url{http://dx.doi.org/10.7153/mia-15-86}.

\bibitem{Seiffert-Jiang-Qi.tex}
W.-D. Jiang, J. Cao, and F. Qi, \textit{Two sharp inequalities for bounding the Seiffert mean by the arithmetic, centroidal, and contra-harmonic means}, available online at \url{http://arxiv.org/abs/1201.6432}.

\bibitem{neuman-sandor-mean.tex}
W.-H. Li and F. Qi, \textit{A unified proof of inequalities and some new inequalities involving Neuman-S\'andor mean}, available online at \url{http://arxiv.org/abs/1312.3500}.

\bibitem{Toader-M-Li-Zheng-Rev.tex}
W.-H. Li and M.-M. Zheng, \textit{Some inequalities for bounding Toader mean}, J. Funct. Spaces Appl. \textbf{2013} (2013), Article ID~394194, 5~pages; Available online at \url{http://dx.doi.org/10.1155/2013/394194}.

\bibitem{Hua-Qi-toader310}
S.-L. Qiu and J.-M. Shen, \emph{On two problems concerning means}, Journal of Hangzhou Insitute of Electronic Engineering \textbf{17} (1997), no.~3, 1\nobreakdash--7. (Chinese)

\bibitem{Hua-Qi-toader31}
Gh. Toader, \emph{Some mean values related to the arithmetic-geometric mean}, J. Math. Anal. Appl. \textbf{218} (1998), no.~2, 358\nobreakdash--368; \url{http://dx.doi.org/10.1006/jmaa.1997.5766}.

\bibitem{Hua-Qi-toader39}
M. Vuorinen, \emph{Hypergeometric functions in geometric function theory}, in: Special Functions and Differential Equations, Proceedings of a Workshop held at The Institute of Mathematical Sciences, Madras, India, January 13-24, 1997, Allied Publ., New Delhi, 1998, 119\nobreakdash--126.

\bibitem{Hua-Qi-toader32}
M.-K. Wang, Y.-M. Chu, S.-L. Qiu, and Y.-P. Jiang, \emph{Bounds for the perimeter of an ellipse}, J. Approx. Theory \textbf{164} (2012), no.~7, 928\nobreakdash--937; Availble online at \url{http://dx.doi.org/10.1016/j.jat.2012.03.011}.

\end{thebibliography}

\documentclass[reqno]{amsart}
\usepackage{amssymb,setspace}
\usepackage{ifpdf}
\ifpdf
 \usepackage[hyperindex,pagebackref]{hyperref}%
\else
 \expandafter\ifx\csname dvipdfm\endcsname\relax
 \usepackage[hypertex,hyperindex,pagebackref]{hyperref}%
 \else
 \usepackage[dvipdfm,hyperindex,pagebackref]{hyperref}%
 \fi
\fi
\theoremstyle{plain}
\newtheorem{thm}{Theorem}
\numberwithin{equation}{section}
\DeclareMathOperator{\td}{d}
\allowdisplaybreaks[4]

\begin{document}

\title[A double inequality for bounding Toader mean]
{A double inequality for bounding Toader mean by the centroidal mean}

\author[Y.Hua]{Yun Hua}
\address[Hua]{Department of Mathematics, School of Science, Tianjin Polytechnic University, Tianjin City, 300387, China}
\email{\href{mailto: Yun Hua <xxgcxhy@163.com>}{xxgcxhy@163.com}}

\author[F. Qi]{Feng Qi}
\address[Qi]{College of Mathematics, Inner Mongolia University for Nationalities, Tongliao City, Inner Mongolia Autonomous Region, 028043, China}
\email{\href{mailto: F. Qi <qifeng618@gmail.com>}{qifeng618@gmail.com}, \href{mailto: F. Qi <qifeng618@hotmail.com>}{qifeng618@hotmail.com}, \href{mailto: F. Qi <qifeng618@qq.com>}{qifeng618@qq.com}}
\urladdr{\url{http://qifeng618.wordpress.com}}

\begin{abstract}
In the paper, the authors find the best numbers $\alpha$ and $\beta$ such that
$$
\overline{C}\bigl(\alpha a+(1-\alpha)b,\alpha b+(1-\alpha)a\bigr)<T(a,b)
<\overline{C}\bigl(\beta a+(1-\beta)b,\beta b+(1-\beta)a\bigr)
$$
for all $a,b>0$ with $a\ne b$, where $\overline{C}(a,b)={2\bigl(a^2+ab+b^2\bigr)}{3(a+b)}$ and $T(a,b)=\frac{2}{\pi}\int_{0}^{{\pi}/{2}}\sqrt{a^2{\cos^2{\theta}}+b^2{\sin^2{\theta}}}\,\td\theta$
denote respectively the centroidal mean and Toader mean of two positive numbers $a$ and $b$.
\end{abstract}

\subjclass[2010]{Primary 26E60; Secondary 26D20, 33E05}

\keywords{Toader mean; centroidal mean; complete elliptic integral; double inequality}

\thanks{This paper was typeset using \AmS-\LaTeX}

\maketitle

\section{Introduction}

In~\cite{Hua-Qi-toader31}, Toader introduced a mean
\begin{align}\label{eq1.1}
T(a,b)&=\dfrac{2}{\pi}\int_{0}^{\pi/2}\sqrt{a^2\cos^2\theta+b^2\sin^2\theta}\,\td\theta\\
&=
\begin{cases}
\dfrac{2a}\pi\mathcal{E}\Biggl(\sqrt{1-\biggl(\dfrac{b}a\biggr)^2}\,\Biggr),&a>b,\\
\dfrac{2b}\pi\mathcal{E}\Biggl(\sqrt{1-\biggl(\dfrac{a}{b}\biggr)^2}\,\Biggr),&a<b,\\
a,&a=b,
\end{cases}
\end{align}
where
$$
{\mathcal{E}}={\mathcal{E}}(r)=\int_{0}^{\pi/2}\sqrt{1-r^2\sin^2\theta}\,\td\theta
$$
for $r\in[0,1]$ is the complete elliptic integral of the second kind.
\par
In recent years, there have been plenty of literature, such as~\cite{Hua-Qi-toader35, Hua-Qi-toader33, Hua-Qi-toader34, Jiang-toader2.tex, jiang-qi-toader-bound.tex, Jiang-Neuman.tex, jiang-qi-October.tex, background-Jiang-Qi.tex, background-Jiang-Qi-revised.tex, Seiffert-Jiang-Qi.tex, neuman-sandor-mean.tex, Toader-M-Li-Zheng-Rev.tex, Hua-Qi-toader32}, dedicated to bounding Neuman-S\'andor's, Seiffert's, Toader's, and other means which are related to complete elliptic integrals of the second kind.
\par
For $p\in \mathbb{R}$ and $a,b>0$, the centroidal mean $\overline{C}(a,b)$ and the $p$-th power mean $M_{p}(a,b)$ are defined respectively by
\begin{equation}\label{eq1.2}
 \overline{ C}(a,b)=\frac{2\bigl(a^2+ab+b^2\bigr)}{3(a+b)}
\end{equation}
and
\begin{equation}\label{eq1.3}
M_{p}(a,b)=
\begin{cases}
\biggl(\dfrac{a^{p}+a^{p}}{2}\biggr)^{{1}/{p}},&p\ne 0,\\
\sqrt{ab}\,,&p=0.
\end{cases}
\end{equation}
\par
In~\cite{Hua-Qi-toader39}, Vuorinen conjectured that
\begin{equation}\label{eq1.4}
  M_{3/2}(a,b)<T(a,b)
\end{equation}
for all $a,b>0$ with $a\ne b$. This conjecture was verified by Qiu and Shen~\cite{Hua-Qi-toader310} and by Barnard, Pearce, and Richards~\cite{Hua-Qi-toader311}.
In~\cite{Hua-Qi-toader312}, Alzer and Qiu presented that
\begin{align}\label{eq1.5}
  T(a,b)<M_{(\ln2)/\ln(\pi/2)}(a,b)
\end{align}
for all $a,b>0$ with $a\ne b$, which gives a best possible upper bound for Toader mean in terms of the power mean.
\par
Very recently, Chu, Wang, and Ma proved in~\cite{Hua-Qi-toader313} that the double inequality
\begin{equation}\label{eq1.6}
C\bigl(\alpha a+(1-\alpha)b,\alpha b+(1-\alpha)a\bigr)<T(a,b)<C\bigl(\beta a+(1-\beta)b,\beta b+(1-\beta)a\bigr)
\end{equation}
is valid for all $a,b>0$ with $a\ne b$ if and only if $\alpha\le \frac34$ and $\beta\ge \frac12+\frac{\sqrt{4\pi-\pi^2}\,}{2\pi}$, where $C(a,b)=\frac{a^2+b^2}{a+b}$ is the contraharmonic mean.
\par
For positive numbers $a, b>0$ with $a\ne b$, let
\begin{equation}\label{J(x)-dfn-eq}
J(x)=\overline{C}\bigl(xa+(1-x)b,xb+(1-x)a\bigr)
\end{equation}
on $\bigl[\frac{1}{2},1\bigr]$. It is easy to see that $J(x)$ is continuous and strictly increasing on $\bigl[\frac{1}{2},1\bigr]$.
Now it is much natural to ask a question: What are the best constants $\alpha\ge\frac12$ and $\beta\le1$ such that the double inequality
\begin{equation}\label{seiffert-th1.1-constant}
\overline{C}\bigl(\alpha a+(1-\alpha)b,\alpha b+(1-\alpha)a\bigr)<T(a,b)
<\overline{C}\bigl(\beta a+(1-\beta)b,\beta b+(1-\beta)a\bigr)
\end{equation}
holds for $a,b>0$ with $a\ne b$? This problem can be affirmatively answered by the following theorem which is the main result of this paper.

\begin{thm}\label{th1.1}
For positive numbers $a,b>0$ with $a\ne b$, the double inequality~\eqref{seiffert-th1.1-constant} is valid if and only if $\alpha\le \frac{1}{2}\bigl(1+\frac{\sqrt{3}\,}{2}\bigr)$ and $\beta\ge \frac{1}{2}+\frac{1}{2}\sqrt{\frac{12}{\pi}-3}\,$.
\end{thm}

\section{Proof of Theorem~\ref{th1.1}}

For $0<r<1$, denote $r'=\sqrt{1-r^2}\,$. It is known that Legendre's complete elliptic integrals of the first and second kind are defined respectively by
$$
\begin{cases}\displaystyle
{\mathcal{K}}={\mathcal{K}}(r)=\int_{0}^{{\pi}/{2}}\frac1{\sqrt{1-r^2\sin^2\theta}}\td \theta,\\
{\mathcal{K}}'={\mathcal{K}}'(r)={\mathcal{K}}(r'),\\
{\mathcal{K}}(0)=\dfrac\pi2,\\
{\mathcal{K}}(1^-)=\infty
\end{cases}
$$
and
$$
\begin{cases}\displaystyle
{\mathcal{E}}={\mathcal{E}}(r)=\int_{0}^{\pi/2}\sqrt{1-r^2\sin^2\theta}\,\td\theta,\\
{\mathcal{E}}'={\mathcal{E}}'(r)={\mathcal{E}}(r'),\\
{\mathcal{E}}(0)=\dfrac\pi2,\\
{\mathcal{E}}(1^-)=1.
\end{cases}
$$
See~\cite{Hua-Qi-toader36, Hua-Qi-toader37}.
For $0<r<1$, the following formulas were presented in~\cite[Appendix~E, pp.~474\nobreakdash--475]{Hua-Qi-toader38}:
\begin{gather*}
 \frac{\td \mathcal{K}}{\td r}=\frac{\mathcal{E}-(r')^2{\mathcal{K}}}{r(r')^2},\quad
\frac{\td \mathcal{E}}{\td r}=\frac{\mathcal{E}-\mathcal{K}}{r},\quad
\frac{\td ({\mathcal{E}}-(r')^2{\mathcal{K}})}{\td r}=r{\mathcal{K}},\\
\frac{\td ({\mathcal{K}}-{\mathcal{E}})}{\td r}=\frac{r{\mathcal{E}}}{(r')^2},
\quad
{\mathcal{E}}\biggl(\frac{2\sqrt{r}}{1+r}\biggr)=\frac{2\mathcal{E}-(r')^2\mathcal{K}}{1+r}.
\end{gather*}
\par
For simplicity, denote
\begin{equation*}
\lambda=\frac{1}{2}\biggl(1+\frac{\sqrt{3}\,}{2}\biggr)\quad \text{and}\quad \mu=\frac{1}{2}+\frac{1}{2}\sqrt{\frac{12}{\pi}-3}\,.
\end{equation*}
It is clear that, in order to prove the double inequality~\eqref{seiffert-th1.1-constant}, it suffices to show
\begin{equation}\label{eq2.1}
  T(a,b)>\overline{C}\bigl(\lambda a+(1-\lambda)b,\lambda b+(1-\lambda)a\bigr)
\end{equation}
and
\begin{equation}\label{eq2.2}
  T(a,b)<\overline{C}\bigl(\mu a+(1-\mu)b,\mu b+(1-\mu)a\bigr).
\end{equation}
From~\eqref{eq1.1} and~\eqref{eq1.2} we see that both $T(a,b)$ and $\overline{C}(a,b)$ are symmetric and homogenous of degree $1$. Hence, without loss of generality, we assume that $a>b$. Let $t=\frac{b}a\in(0,1)$ and $r=\frac{1-t}{1+t}\in(0,1)$ and let $p\in \bigl(\frac12,1\bigr)$. Then
\begin{multline}\label{eq2.3}
\begin{aligned}
&\quad T(a,b)-\overline{C}\bigl(pa+(1-p)b,pb+(1-p)a\bigr)\\
&=\frac{2a}{\pi}\mathcal{E}\Biggl(\sqrt{1-\biggl(\frac{b}{a}\biggr)^2}\,\Biggr)
\end{aligned}\\
-2a\frac{[p+(1-p)b/a]^2+[p+(1-p)b/a](pb/a+1-p)+(pb/a+1-p)^2}{3(1+b/a)}\\
=\frac{2a}{\pi}\mathcal{E}\bigl(\sqrt{1-t^2}\,\bigr)-2a\frac{[p+(1-p)t]^2 +[p+(1-p)t](pt+1-p)+(pt+1-p)^2}{3(1+t)}\\
\begin{aligned}
&=\frac{2a}{\pi}\frac{2\mathcal{E}-\bigl(1-r^2\bigr)\mathcal{K}}{1+r}-a\frac{(1-2p)^2r^2+3}{3(1+r)}\\
&=\frac{a}{1+r}\biggl\{\frac{2}{\pi}\bigl[2\mathcal{E}-\bigl(1-r^2\bigr)\mathcal{K}\bigr] -\frac{1}{3}(1-2p)^2r^2-1\biggr\}.
\end{aligned}
\end{multline}
Let
\begin{align}\label{eq2.4}
f(r)=\frac{2}{\pi}\bigl[2\mathcal{E}-\bigl(1-r^2\bigr)\mathcal{K}\bigr]-\frac{1}{3}(1-2p)^2r^2-1,
\end{align}
and let $f_1(r)=rf'(r)$ and $f_2(r)=\frac{f_1'(r)}r$. Then, by standard argument, we have
\begin{gather*}
  f(0)=0,\quad f_1(0)=0,\quad f_2(0)=1-\frac{4}{3}(1-2p)^2,\\
  f(1^-)=\frac{4}{\pi}-1-\frac{1}{3}(1-2p)^2,\quad
 f_1(1^-)=\frac{2}{\pi}-\frac{2}{3}(1-2p)^2,\quad   f_2(1^-)=+\infty,\\
f_1(r)=\frac{2}{\pi}\bigl[\mathcal{E}-\bigl(1-r^2\bigr)\mathcal{K}\bigr]-\frac{2}{3}(1-2p)^2r^2,\quad
   f_2(r)=\frac{2}{\pi}\mathcal{K}-\frac{4}{3}(1-2p)^2,
\end{gather*}
\par
When $p=\lambda=\frac{1}{2}\bigl(1+\frac{\sqrt{3}\,}{2}\bigr)$, it follows that $f_2(0)=0$. An easy argument leads to $f(r)>0$ for $r\in(0,1)$. Together with this, the inequality~\eqref{eq2.1} follows from~\eqref{eq2.3} and~\eqref{eq2.4}.
\par
When $p=\mu=\frac{1}{2}+\frac{1}{2}\sqrt{\frac{12}{\pi}-3}\,$, it is simple to derive that
$$
f(1^-)=0,\quad f_1(1^-)=\frac{2(\pi-3)}{\pi}>0,\quad f_2(0)=\frac{5\pi-16}{\pi}<0.
$$
Consequently, considering the monotonicity of $f_2(r)$, it is deduced that there exists $r_0\in(0,1)$ such that $f_2(r)<0$ on $(0,r_0)$ and $f_2(r)>0$ on $(r_0,1)$. Hence, the function $f_1(r)$ is strictly decreasing on $(0,r_0)$ and strictly increasing on $(r_0,1)$.
Similarly, there exists $r_1\in(0,1)$ such that $f_1(r)<0$ on $(0,r_1)$ and $f_1(r)>0$ on $(r_1,1)$. Thus, the function $f(r)$ is strictly decreasing on $(0,r_1)$ and strictly increasing on $(r_1,1)$.
As a result, the inequality~\eqref{eq2.2} follows.
\par
If $p>\lambda$, then $f_2(r)<0$. From the continuity of $f(r)$, $f_1(r)$, and $f_2(r)$, it follows that there exists $\delta_1=\delta_1(p)>0$ such that $f(r)<0$ on $(0,\delta_1)$. Combining this with~\eqref{eq2.3} and~\eqref{eq2.4} yields that $T(a,b)<\overline{C}\bigl(p a+(1-p)b,p b+(1-p)a\bigr)$ for $\frac{b}a\in\bigl(\frac{1-\delta_1}{1+\delta_1},1\bigr)$.
If $p<\mu$, then $f(1^-)>0$. Hence, there exists $\delta_2=\delta_2(p)\in(0,1)$ such that $f(r)>0$ on $(1-\delta_2,1)$. Combining this with~\eqref{eq2.3} and~\eqref{eq2.4} reveals that $T(a,b)>\overline{C}\bigl(p a+(1-p)b,p b+(1-p)a\bigr)$ for $\frac{b}a\in(0,\delta_2/(2-\delta_2))$. These imply that the constants $\lambda$ and $\mu$ are the best possible.
The proof of Theorem~\ref{th1.1} is complete.

\subsection*{Acknowledgements}
The authors thank the anonymous referee for his/her careful reading and helpful corrections to the original version of this paper.
\par
The first author was partially supported by the Project of Shandong Province Higher Educational Science and Technology Program under Grant No. J11LA57, China.

\end{document}